The Goodyear blimp flies at a cruising speed of 16 meters per second. The blimp requires 500 meters to go from being parked to being at cruising speed. What is the acceleration of the blimp?
1
Expert's answer
2019-10-09T10:42:40-0400

We need to use:


"v_f^2-v_i^2=2as"

"16^2-0^2=2(500)a"

"a=0.256\frac{m}{s^2}"